An investment of $100 today will avoid $100,000 of environmental damage in 100 years. Calculate net present value to answer each of the following: a. At a discount rate of 10%, is this investment a good idea? b. At a discount rate of 1%, is this investment a good idea?
Answer #1

t = 100 y , FV = $ 100000

a) NPV = 100000/(1 + 10%)100 - 100 = - $ 92.74

NPV is negative, investment not a good idea.

b) NPV = 100000/(1 + 1%)100 - 100 = $ 36871.12

NPV is positive, investment is a good idea.